OpenStudy (anonymous):

I will give metalThe area of the circle above is 1,661.06 units2. What is the circumference of the circle? (Use 3.14 for .) A. 72.22 units B. 830.53 units C. 553.69 units D. 144.44 units

OpenStudy (anonymous):

Formula for area of circle is \[\pi*r^{2}\] so \[\pi*r^{2}=1661.06\] so \[1661.06\div \pi = r^{2}\] \[r =\sqrt{1661.06 \div \pi} = 23\] so the radius is 23 units. Circumference is equal to 2*pi*r therefore circumference is about 144.44

OpenStudy (anonymous):

its D